OpenStudy (anonymous):

Lydia received a parking ticket because she parked ten feet from a fire hydrant. She had estimated her distance from the fire hydrant as being sixteen feet. What was her percent error? sixteen percent thirty-eight percent sixty percent sixty-three percent

OpenStudy (anonymous):

i belive it would be 36%

OpenStudy (anonymous):

Thanks Cram.

OpenStudy (anonymous):

Very simple ..The formula for this is: \[Increase/Decrease \times 100\] Increase = N - Original=16 - 10 = 6 Therefore:: \[\frac{ 6 }{ 10 } \times 100 = 60 %\]

OpenStudy (anonymous):

Thanks Adjax

OpenStudy (anonymous):

which answer is right 36% or 60%

OpenStudy (anonymous):

@_lilb_to_smooth_4_U_

OpenStudy (anonymous):

60% @Adjax

OpenStudy (anonymous):

yeah :thumps_up: